OpenStudy (anonymous):

Please explain step by step. Let f(x)= x/x^2+1 use the first derivative test to determine which critical numbers if any give relative extrema determine the intervals on which the function is increasing or decreasing by evaluating f'(x) at appropriate test values. Based on the sign of the derivative changes between intervals determine where the function has relative extrema .

OpenStudy (solomonzelman):

Did you get the \(f'(x)\) ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

There are three cases when critical numbers occur. \(\sf\color{red}{■ ~~~Case ~1}\) If any closed-x-interval boundaries are given then, these closed boundaries are critical numbers. \(\sf\color{red}{■ ~~~Case ~2}\) Any x-values that satisfy \(f'(x)=0\) are critical numbers. (Well, if they are part of function's domain) \(\sf\color{red}{■ ~~~Case ~3}\) Any x-values at which \(f'(x)\) is undefined. Provided that at these x-values the \(f(x)\) is defined.

OpenStudy (solomonzelman):

For the derivative, [the typical thing here is to] use the "Quotient Rule"

OpenStudy (solomonzelman):

What was your derivative, can you show me?

OpenStudy (anonymous):

1/(x^2 +1) - 2x^2/(x^2+1)^2

OpenStudy (anonymous):

But that can be simplified

OpenStudy (solomonzelman):

that is rather likely, i mean the way you wrote, if you did, \(\color{#000000 }{ \displaystyle \frac{d }{dx} \left[x(x^2+1)^{-1}\right]=(x)'(x^2+1)^{-1}+x\left((x^2+1)^{-1}\right)' }\) \(\color{#000000 }{ \displaystyle \frac{d }{dx} \left[x(x^2+1)^{-1}\right]=1\cdot (x^2+1)^{-1}+x\left(-(x^2+1)^{-2}\times 2x\right) }\) \(\color{#000000 }{ \displaystyle \frac{d }{dx} \left[x(x^2+1)^{-1}\right]=\frac{1}{x^2+1}+\frac{-2x^2}{(x^2+1)^2} }\) Good job

OpenStudy (solomonzelman):

is this how you did it? (through product rule, with negative exponent) I like that better than quotient, btw.

OpenStudy (solomonzelman):

In any case, the derivative is right so let's move on.

OpenStudy (anonymous):

Ok

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f'(x)=\frac{1}{x^2+1}-\frac{ 2x^2 }{(x^2+1)^2} }\) but, I think we would need to combine anyway \(\color{#000000 }{ \displaystyle f'(x)=\frac{x^2+1}{(x^2+1)^2}-\frac{ 2x^2 }{(x^2+1)^2} }\) \(\color{#000000 }{ \displaystyle f'(x)=\frac{x^2+1- 2x^2 }{(x^2+1)^2} }\) that is simpler looking... \(\color{#000000 }{ \displaystyle f'(x)=\frac{1-x^2 }{(x^2+1)^2} }\)

OpenStudy (solomonzelman):

You are NOT given any closed-interval boundaries, so case 1 (for critical numbers) is excluded. Case 2 is also excluded, since \(x^2+1\) is defined for all real x.

OpenStudy (solomonzelman):

So, only remains case 3.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0=\frac{1-x^2 }{(x^2+1)^2} }\) please solve for x.

OpenStudy (anonymous):

One moment

OpenStudy (solomonzelman):

it will be quick if you multiply both sides times \((x^2+1)^2\)

OpenStudy (anonymous):

What.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0=\frac{1-x^2 }{(x^2+1)^2} }\) What I mean is, \(\color{#000000 }{ \displaystyle 0~\color{red}{\times (x^2+1)^2}=\frac{1-x^2 }{(x^2+1)^2}~\color{red}{\times (x^2+1)^2} }\)

OpenStudy (anonymous):

Isn't it just one?

OpenStudy (solomonzelman):

For the left side you know that anything times 0 is equal to ..... For the right side, \((x^2+1)^2\) will cancel out.

OpenStudy (solomonzelman):

x=1, is a solution, but not the only.

OpenStudy (anonymous):

Ok how are their other solutions?

OpenStudy (solomonzelman):

yes, there is another solution besides x=1.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0=\frac{1-x^2 }{(x^2+1)^2} }\) you should be capable of solving this. (You got the derivative. This is so much simpler)

OpenStudy (anonymous):

Negative 1?

OpenStudy (solomonzelman):

yes:)

OpenStudy (anonymous):

I was out because I was in the hospital when we were taught how to do these. I have a very one track mind.

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Thank you :)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0=\frac{1-x^2 }{(x^2+1)^2} }\) \(\color{#000000 }{ \displaystyle 0~\color{red}{\times (x^2+1)^2}=\frac{1-x^2 }{(x^2+1)^2}~\color{red}{\times (x^2+1)^2} }\) \(\color{#000000 }{ \displaystyle 0=\frac{1-x^2 }{\cancel{(x^2+1)^2}}~\color{}{\times \cancel{(x^2+1)^2}} }\) \(\color{#000000 }{ \displaystyle 0=1-x^2 }\) \(\color{#000000 }{ \displaystyle x^2=1 }\) (Will do it precisely, instead of making the \(\pm\) pop from nowhere) I will take the square root of both sides, \(\color{#000000 }{ \displaystyle \sqrt{x^2}=\sqrt{1} }\) \(\color{#000000 }{ \displaystyle |x|=1 }\) \(\color{#000000 }{ \displaystyle x=\pm 1 }\)

OpenStudy (solomonzelman):

(That is the definition of absolute value \(|x|=\sqrt{x^2}\) (for real numbers))

OpenStudy (solomonzelman):

So, your critical numbers are?

OpenStudy (anonymous):

The absolute value of 1?

OpenStudy (anonymous):

Or 1 and -1

OpenStudy (solomonzelman):

yes, the critical numbers are;\ x=1 x=-1

OpenStudy (solomonzelman):

And now, plug these numbers into the function and tell me what you get for each one of them.

OpenStudy (anonymous):

1/2 and -1/2

OpenStudy (solomonzelman):

Yes, very good.

OpenStudy (solomonzelman):

In fact, f(1)=1/2 is absolute maximum (max of y=1/2 that occurs at x=1) f(-1)=-1/2 is absolute minimumn (min of y=-1.2 that occurs at x=-1) (if I find a need, we will refer to this statement later to show it)

OpenStudy (solomonzelman):

When the slope is negative the function is decreasing. When the slope is positive the function is increasing. Am I right?

OpenStudy (anonymous):

I believe so XD

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

The derivative is the [instantaneous] slope. (This is by defintion, and if you want, I can precisely explain why this is so after we are done with this problem)

OpenStudy (solomonzelman):

So, equivalently correct statements would be, When the \(\color{#000000 }{ \sf derivative }\) is negative the function is decreasing. When the \(\color{#000000 }{ \sf derivative }\) is positive the function is increasing.

OpenStudy (solomonzelman):

(If you ever lose me, or if I am going too fat/too slow then let me know)

OpenStudy (anonymous):

No, you're good

OpenStudy (solomonzelman):

Ok, and we actually have the derivative expression/function. So, we are able to figure when the slope/derivative is negative and when is positive. (you know that positive means greater than 0, and negative means less than zero)

OpenStudy (solomonzelman):

If you recall, I have simplified the derivative to be (if yo don't recall, then simplifying this is fairly easy) \(\color{#000000 }{ \displaystyle f'(x)=\frac{1-x^2 }{(x^2+1)^2} }\)

OpenStudy (solomonzelman):

So for negative slopes \(\color{#000000 }{ \displaystyle 0>\frac{1-x^2 }{(x^2+1)^2} }\) and for positive slopes \(\color{#000000 }{ \displaystyle 0<\frac{1-x^2 }{(x^2+1)^2} }\)

OpenStudy (solomonzelman):

now, you need to solve and tell me where the slope/derivative is negative and where [the derivative] is positive.

OpenStudy (anonymous):

So when x is equal to or less than 1 it's negative? And when x is equal to or greater than 2 it's positive?

OpenStudy (solomonzelman):

well, no...

OpenStudy (solomonzelman):

Negative slope, \(\color{#000000 }{ \displaystyle 0>\frac{1-x^2 }{(x^2+1)^2} }\) \(\color{#000000 }{ \displaystyle 0>1-x^2 }\) \(\color{#000000 }{ \displaystyle x^2>1 }\) \(\color{#000000 }{ \displaystyle \sqrt{x^2}>\sqrt{1} }\)

OpenStudy (solomonzelman):

keep going from there please.

OpenStudy (anonymous):

I honestly don't know what I'm supposed to do there

OpenStudy (solomonzelman):

simplify that and solve for x.

OpenStudy (anonymous):

I'm barley passing calculus and I barely even passed the past two years of math.

OpenStudy (solomonzelman):

Then, I guess you have to review the previous maths untill you can answer any of these questions. You can't move on like this.

OpenStudy (anonymous):

It's my last year of high school. I get through this and I can go back to algebra in collage

OpenStudy (solomonzelman):

but, how can you do calculus without knowing algebra?

OpenStudy (solomonzelman):

(btw, I have recently seen a user with a similar problem, and I think I motivated this user to review algebra)

OpenStudy (anonymous):

I'd like to pass on the whole motivation thing. I had awful teachers until I changed schools and now my teachers are just trying to keep me afloat. I'm an A+ student.... But in math I'm more like a D student

OpenStudy (anonymous):

It's hard enough without your judgement

OpenStudy (solomonzelman):

Just look the common course outline for algebra and try to nail all of the things possible.

OpenStudy (solomonzelman):

have a good look!

OpenStudy (anonymous):

thank you for your time.

OpenStudy (solomonzelman):

Anytime.